# STORM: Segment, Track, and Object Re-Localization from a Single 3D Model

### v1

@@ Abstract
Accurate 6D pose estimation and tracking are fundamental capabilities for physical AI systems such as robots. However, existing approaches typically rely on a manually annotated segmentation mask of the target in the first frame, which is labor-intensive and leads to reduced performance when faced with occlusions or rapid movement. To address these limitations, we propose STORM (Segment, Track, and Object Re-localization from a single 3D Model), an open-source robust real-time 6D pose estimation system that requires no manual annotation. STORM employs a novel three-stage pipeline combining vision-language understanding with self-supervised feature matching: contextual object descriptions guide localization, self-cross-attention mechanisms identify candidate regions, and a segmentation model produces precise masks for accurate pose estimation. Another key innovation is our automatic re-registration mechanism that detects tracking failures through feature similarity monitoring and recovers from severe occlusions or rapid motion. STORM achieves state-of-the-art accuracy on challenging industrial datasets featuring multi-object occlusions, high-speed motion, and varying illumination, while operating at real-time speeds without additional training. This annotation-free approach significantly reduces deployment overhead, providing a practical solution for modern applications, such as flexible manufacturing and intelligent quality control. Code: https://github.com/dengyufuqin/Storm
@@ 1 Introduction
Estimating the rigid six-degree-of-freedom (6D) transformation between an object and the camera—also known as object pose estimation—is a critical task in robotic manipulation (Kappler et al. 2018; Wen et al. 2022) and augmented reality (Marchand, Uchiyama, and Spindler 2015). Despite recent significant advances (He et al. 2020, 2021; Labbé et al. 2020; Park, Patten, and Vincze 2019; Wen et al. 2020), current methods face critical limitations that restrict their deployment in real-world scenarios. First, most methods operate at the instance level, failing to generalize to novel objects. Second, they heavily depend on prior information such as object masks or bounding boxes (Wen et al. 2024; Hu et al. 2019; Do et al. 2018), limiting their applicability. These limitations stem from their reliance on extensive reference information as model input, creating a bottleneck for practical deployment where systems must handle diverse, previously unseen objects.
Recent advances partially address these limitations through reference-based approaches. CNOS (Nguyen et al. 2023) and PerSAM (Zhang et al. 2023) generate object masks in query frames by either rendering 3D models into reference images or using user-provided reference images. These methods employ pre-trained vision models (e.g., DINOv2 (Oquab et al. 2023)) to extract visual features and identify target objects through similarity matching. However, their performance degrades in cluttered scenes with occlusions due to their simplistic matching mechanism (e.g., cosine similarity). Moreover, they first segment all objects before performing individual comparisons, incurring substantial computational overhead. This makes it challenging to maintain object consistency across time.
@@
To address these challenges, we propose STORM (Segment, Track, and Object Re-localization from a single 3D Model), a unified framework for 6D object pose estimation that operates with minimal priors. Using only a 3D model, STORM generates accurate object masks and recovers 6D poses, while performing object re-localization through a dedicated memory pool to enable reliable, continuous tracking. STORM comprises three key stages:First, we render the 3D model into 16 spherically distributed reference views and use a Large Language Model (LLM) to generate semantic descriptions compatible with CLIP (Radford et al. 2021). Second, we extract spatial features from the reference and query images using object detection models such as DINOv2. To bridge the domain gap between rendered references and real-world imagery, we introduce Hierarchical Spatial Fusion Attention (HSFA)—a mechanism that fuses multi-view self-attention and cross-attention. Third, we convert the resulting regions into effective prompts (centroids and bounding boxes) for foundation segmentation models such as FastSAM (Zhao et al. 2023) and SAM2 (Ravi et al. 2024), yielding precise object masks. To ensure robust tracking, we introduce a lightweight tracking-loss classifier that penalizes tracking failures using an efficient object memory mechanism. As a result, STORM automatically re-registers lost objects, achieving robust and reliable object tracking even under occlusion and rapid motion. Overall, we make the following contributions:
  • We propose STORM, a unified framework for robust object pose estimation and tracking using only a single 3D model that renders multi-view reference images for accurate pose predictions.
@@
  • To evaluate the tracking robustness, we create a new dataset, that represents object tracking as a classification task, by comparing the input image to a reference image. Our dataset covers diverse track-loss scenarios, including occlusion and rapid movement.
@@
Figure 2: Overview of STORM, composed of two subsystems: the Segmenting Object Module (SOM) and the Tracking Object Module (TOM). SOM leverages semantic and spatial information from reference images or multi-view images. By integrating both intra-image and inter-image attention modules, it captures spatial cues of the query frame through a local and a global attention block, to produce a segmented mask. TOM classifies the output of the tracking module and utilizes the memory of the SOM module to perform re-registration when tracking fails, thereby ensuring robust and continuous tracking performance.
@@ 2 Background and Related Work
STORM is at the intersection of reference-based object segmentation and object pose tracking.
@@ 3 STORM
We propose STORM that addresses two key limitations of prior approaches: (i) the reliance on an initial object mask for pose estimation, and (ii) the lack of mechanisms to recover from tracking failures. STORM leverages efficient, pre-established models for object segmentation, language understanding, and pose estimation. As a result, it provides a fully automated pipeline: given a single 3D model, STORM localizes and segments the target object, infers its semantic attributes, estimates its 6D pose, and performs self-correcting tracking across frames. The inference process in STORM comprises two main components: the segmenting object module (SOM) and tracking object module (TOM). Let us now describe each component in detail.
Object Segementation from a Single 3D Model
STORM robustly performs object segmentation by utilizing the provided 3D model in conjunction with a novel attention-driven feature matching strategy.
@@
First, STORM transforms the 3D model into a set of reference images to guide the subsequent segmentation process, as depicted in Figure 2 1. To capture the object’s full spatial characteristics, the 3D model (M) is rendered into 16 textured views ($I_{k}$) under rotations ($R_{k}\in\mathrm{SO}(3)$) that uniformly sample the unit sphere (Li et al. 2024). These multi-view renderings enable the system to learn the object’s appearance from all angles, facilitating robust cross-view matching with minimal computational overhead.
@@
To acquire semantic information, we feed the frontal view $I_{\mathrm{front}}$ and a short prompt $p$ into a vision language model (gemma-3n-E4B-it) (Team 2025), which produces a concise description of the object $T$ (avoiding redundant information). As depicted in Figure 2 2, the image is encoded by CLIP into a vector $e_{t}\in\mathbf{R}^{d}$. Rather than altering the backbone’s core features, we inject $e_{t}$ into the downstream feature map $F\in\mathbf{R}^{H\times W\times C}$ by modifying its channel-wise normalization parameters. Specifically, we compute semantic offsets
@@
To bridge the distribution gap between rendered and the real-world images and to capture the spatial information of the object, we introduce Hierarchical Spatial Fusion Attention (HSFA). As shown in Figure 2 3, HSFA uses the feature representations from both the real and rendered frames. It fuses the distinct contexts and generates prompts for the subsequent segmentation module, leading to high-quality mask predictions. Overall, HSFA processes the input as follows:
@@
    Capturing inter-view relationships and rich spatial representations. Patch embeddings from each multi-view image are first processed independently using self-attention. These per-image embeddings are then concatenated across all views and passed through a joint local attention module, allowing the model to capture inter-view dependencies.
@@
    Extracting salient query representations. The query frame’s patch embeddings are processed via self-attention. To preserve the original DinoV2 features, HSFA performs the first round of cross-attention between the query embeddings and the raw multi-view embeddings (i.e., before spatial-attention refinement). In subsequent cross-attention layers, the query frame attends to the outputs of the previous spatial-attention block, promoting faster convergence and enhancing object salience.
@@
    Composing joint representations. The Hierarchical Spatial Fusion Attention block integrates the query and multi-view embeddings, as described above. This fusion block is repeated $n$ times to iteratively refine the joint features. Finally, an upsampling convolutional block generates a region-of-interest for object localization.
@@
The full algorithm is available in the Appendix. The output of the HSFA module is a set of fused feature representations that integrate spatial cues from the query frame and multi-view inputs. From these features, the module derives a structured set of prompts—specifically, representative point prompts (including the centroid and subregion centroids) and a tight bounding box around the region of interest—which are used to guide the segmentation model.
@@
where $\mathrm{TP}$ is the number of true‐positive pixels of the region of interest, $\mathrm{FN}$ is the number of false‐negative pixels, and $\mathrm{FP}$ is the number of false‐positive pixels; $t\in\{0,1\}$ is the ground‐truth label for each pixel, and $p\in[0,1]$ is the predicted probability that the pixel belongs to the region of interest. The hyperparameters are set as $\alpha=0.3$, $\beta=0.7$, $\gamma=2.0$, $\lambda_{T}=1.0$, and $\lambda_{f}=0.2$. This loss formulation strongly penalizes false positives, resulting in a compact, high-precision region of Interest that retains critical edge details for downstream foundation segmentation.
@@
    Initial Registration and Memory Construction. For the first tracking frame, STORM uses its segmentation module to extract an object mask. This mask, along with the object’s 3D model, is passed to the FoundationPose registration module to estimate the object’s 6D pose. At the same time, a cropped RGB patch of the segmented object—referred to as a real crop—is stored in a memory pool. This crop serves as a visual reference for subsequent frames. In each new frame, the tracking module produces a pose estimate, which is compared to the memory pool to detect tracking failures. If the similarity falls below a threshold—due to rapid motion or occlusion—STORM flags the tracker as failed.
@@
    Re-initialization and Tracking Verification. When a failure is detected, STORM automatically re-initializes the tracker by SOM to generate a new segmentation mask and re-estimating the object’s pose. A new real crop from the current frame is also added to the memory pool, ensuring that visual references stay up to date. To enable robust and generalizable comparisons, DINOv2 is used as the feature-extraction backbone. Features from each frame are concatenated and passed through a lightweight tracking-verification subnetwork, which consists of a single self-attention layer followed by a fully connected layer and a sigmoid activation. This network outputs a binary score indicating whether tracking remains valid. The entire mechanism is designed for high efficiency, enabling real-time detection and recovery from tracking failures.
@@ 4 Experimental evaluation
Self-Supervised
STORM (SOM, Ours) 66.1 $\circ$47.1 $\circ$69.1 70.8 89.3 68.5 $0.183$
Supervised
ZebraPoseSAT–EffNetB4 $\circ$51.6 72.1 71.8 $\circ$68.9 $\circ$73.1 $\circ$67.5 0.080
Mask R-CNN 37.5 54.4 48.9 47.1 42.9 46.2 $\circ$0.100
Unsupervised
@@
Table 2: STORM (SOM) achieves high-quality segmentation via a self-supervised approach. Comparison of segmentation performance across supervision levels. mAP scores are reported across five BOP datasets. STORM (SOM) achieves the highest mean mAP among self-supervised, supervised and unsupervised methods, while also being highly efficient in runtime. Best values are in bold, and $\circ$ depicts the second best. SOM stands for the Segmenting Object Module in the STORM architecture.
@@
Setup
We train the SOM module using the sub‐datasets provided by the BOP Challenge. We adopted a three‐layer HSFA architecture and trained the model for approximately $80000$ steps on eight NVIDIA A100 GPUs. For TOM, we employed a custom in‐house dataset (which we open source alongside this publication) and trained the network for $15$ epochs on a single NVIDIA A6000 GPU. In both cases, optimization was performed using AdamW (Loshchilov and Hutter 2017). We applied color jittering (Krizhevsky, Sutskever, and Hinton 2012), random rotations, and translations (Simard, Steinkraus, and Platt 2003) as data augmentation strategies. Detailed settings for training and evaluation are provided in Appendix C and  D.
@@
We compare STORM against the strongest publicly available systems on the BOP leaderboard, categorized into supervised (object-specific retraining) and unsupervised (zero-retraining) settings. Supervised baselines include ZebraPoseSAT‑EffiNetB4 (Su et al. 2022), and Mask R-CNN (He et al. 2017), which are fine-tuned per object category. Unsupervised baselines include NOCTIS, LDSeg, MUSE, Prisma‑MPG‑Complex, NIDS, SAM6D (Lin et al. 2024), CNOS and ViewInvDet. Most unsupervised baselines are distributed solely as executable packages via the BOP Challenge website without linked peer-reviewed publications, collectively referred to as the BOP Challenge Leaderboard [^1] https://bop.felk.cvut.cz/challenges/. In contrast, STORM is categorized as a self-supervised method, requiring no manual annotations.
To ensure the proposed method remains entirely label-free, we trained the SOM model on the BOP-Classic-Core training set using rendered reference images from 16 uniformly sampled spherical viewpoints, enabling HSFA to learn object-centric correspondences without any human supervision (see Appendix for complete training details).
Results. Table 2 presents a comparison of supervised and unsupervised segmentation methods across the BOP datasets. Among the supervised baselines, ZebraPoseSAT–EffNetB4 achieves the highest performance, reaching a mean Average Precision (mAP) (Lin et al. 2014) of 67.567.5%. Our zero-shot segmentation module, SOM, achieves a higher mAP of 68.568.5%, with an average inference time of $0.183$ seconds per image. Notably, SOM establishes a new state-of-the-art on four out of the five benchmarks—LM-O, T-LESS, HomebrewedDB, and YCB-Video—without any fine-tuning or retraining for specific object instances. These results underscore SOM’s strong generalization ability.
@@
To answer Q3, we evaluate whether STORM’s tracking module (TOM) can detect when it loses track of an object—that is, whether it exhibits self-awareness of tracking failure. To this end, we synthetically generate a dataset using BOP annotations, enabling TOM to learn to distinguish between successful and failed tracking by leveraging object crops, distractors, and viewpoint variations.
@@
To answer Q4, we compare the performances of different matching strategies on the tracking-loss classification task. Using our tracking dataset, we compare STORM’s tracking module (TOM) with different configurations against a simple cosine similarity baseline. Table 4 provides the accuracy and inference time. Adding a single attention layer significantly improves accuracy from 87.5587.55% (cosine similarity) to 98.3698.36%, with only 1.31.3ms of additional latency per frame. The model without attention already achieves detection accuracy of 95.2395.23% while maintaining the fastest inference at 12.212.2ms. Adding a second attention layer does not improve the accuracy but increasing latency to 16.816.8ms, indicating that a single attention layer optimally balances accuracy and latency. TOM’s attention mechanism is crucial for robust tracking failure detection, substantially outperforming simple feature matching approaches while maintaining inference time suitable for real-time applications.

### v2

# STORM: Segment, Track, and Object Re-Localization from a Single Image

@@ Abstract
Accurate 6D pose estimation and tracking are fundamental capabilities for physical AI systems such as robots. However, existing approaches typically require a pre-defined 3D model of the target and rely on a manually annotated segmentation mask in the first frame, which is labor-intensive and leads to reduced performance when faced with occlusions or rapid movement. To address these limitations, we propose STORM (Segment, Track, and Object Re-localization from a single iMage), an open-source robust real-time 6D pose estimation system that requires no manual annotation. STORM employs a novel three-stage pipeline combining vision-language understanding with feature matching: contextual object descriptions guide localization, self-cross-attention mechanisms identify candidate regions, and produce precise masks and 3D models for accurate pose estimation. Another key innovation is our automatic re-registration mechanism that detects tracking failures through feature similarity monitoring and recovers from severe occlusions or rapid motion. STORM achieves state-of-the-art accuracy on challenging industrial datasets featuring multi-object occlusions, high-speed motion, and varying illumination, while operating at real-time speeds without additional training. This annotation-free approach significantly reduces deployment overhead, providing a practical solution for modern applications, such as flexible manufacturing and intelligent quality control. Code: https://github.com/dengyufuqin/Storm
@@ 1 Introduction
Estimating the rigid six-degree-of-freedom (6D) transformation between an object and the camera—also known as object pose estimation—is a critical task in robotic manipulation (Kappler et al. 2018; Wen et al. 2022) and augmented reality (Marchand, Uchiyama, and Spindler 2015). Despite recent significant advances (He et al. 2020, 2021; Labbé et al. 2020; Park, Patten, and Vincze 2019; Wen et al. 2020), current methods face critical limitations that restrict their deployment in real-world scenarios. First, most methods operate at the instance level, failing to generalize to novel objects. Second, they heavily depend on prior information such as accurate 3D object models, object masks, or bounding boxes (Wen et al. 2024; Hu et al. 2019; Do et al. 2018), limiting their applicability. These limitations stem from their reliance on extensive reference information as model input, creating a bottleneck for practical deployment where systems must handle diverse, previously unseen objects.
Recent advances partially address these limitations through reference-based approaches and object-centric 3D reconstruction techniques.
@@
CNOS (Nguyen et al. 2023) and PerSAM (Zhang et al. 2023) generate object masks in query frames by either rendering 3D models into reference images or using user-provided reference images. These methods employ pre-trained vision models (e.g., DINOv2 (Oquab et al. 2023)) to extract visual features and identify target objects through similarity matching. However, their performance degrades in cluttered scenes with occlusions due to their simplistic matching mechanism (e.g., cosine similarity). Moreover, they first segment all objects before performing individual comparisons, incurring substantial computational overhead. This makes it challenging to maintain object consistency across time.
Recent advances in object-centric 3D reconstruction further complement these reference-based methods. Classical approaches rely on multi-view stereo and RGB-D fusion or SLAM pipelines to build dense scene geometry (Dai et al. ; Dai et al. 2017; Engel, Schöps, and Cremers 2014), but typically operate at the scene level and require long, calibrated sequences. Neural 3D representations, such as object-centric radiance fields (Müller et al. 2022; Cerkezi and Favaro 2024), learn implicit 3D geometry and appearance directly from images and can generalize across object instances, yet they often demand careful optimization and multi-view coverage, limiting their practical use for real-time pose tracking.
@@
To address these challenges, we propose STORM (Segment, Track, and Object Re-localization from a single iMage), a unified framework for 6D object pose estimation that operates with minimal priors. Using only one or more reference images, STORM generates accurate object masks and recovers 6D poses, while performing object re-localization through a dedicated memory pool to enable reliable, continuous tracking. STORM comprises three key stages:First, we take one or more user-provided reference images of the target object and use a Large Language Model (LLM) to generate semantic descriptions compatible with CLIP (Radford et al. 2021). Based on these reference images, SAM3D is then employed to recover the 3D model of the object. Second, we extract spatial features from the reference and query images using visual encoders such as DINOv2. To bridge the domain gap between reference images and real-world imagery, we introduce Hierarchical Spatial Fusion Attention (HSFA)—a mechanism that fuses multi-view self-attention and cross-attention. Third, HSFA produces high-quality segmentation masks of the target object, which are combined with the reconstructed 3D model to estimate its 6D pose. To ensure robust tracking, we introduce a lightweight tracking-loss classifier that penalizes tracking failures using an efficient object memory mechanism. As a result, STORM automatically re-registers lost objects, achieving robust and reliable object tracking even under occlusion and rapid motion. Overall, we make the following contributions:
  • We propose STORM, a unified framework for robust object pose estimation and tracking that requires one or multiple images of the target object to achieve accurate pose predictions.
@@
  • To evaluate the tracking robustness, we create a new dataset that represents object tracking as a classification task, by comparing the input image to a reference image. Our dataset covers diverse track-loss scenarios, including occlusion and rapid movement.
@@
Figure 2: Overview of STORM, which is composed of two subsystems: the Segmenting Object Module (SOM) and the Tracking Object Module (TOM). SOM leverages reference images to generate a 3D model and, using their semantic and spatial information, integrates both intra-image and inter-image attention modules to capture spatial cues of the query frame through local and global attention blocks, producing a segmented mask. TOM classifies the output of the tracking module and utilizes the memory of SOM to perform re-registration when tracking fails, thereby ensuring robust and continuous tracking performance.
@@ 2 Background and Related Work
STORM is at the intersection of reference-based object segmentation, object-centric 3D reconstruction and object pose tracking.
@@
Object-centric 3D Reconstruction. Traditional object-centric 3D reconstruction methods often rely on multi-view geometry and multi-view stereo techniques (Cremers and Kolev 2011), as well as large-scale scene-level reconstruction systems and datasets such as voxel-hashing-based real-time reconstruction (Nießner et al. 2013) and the ScanNet indoor reconstruction benchmark (Dai et al. 2017). These approaches typically require dense, calibrated multi-view imagery or pre-defined CAD models, leading to a trade-off between reconstruction quality, runtime, and deployment cost. More recent work explicitly focuses on object-centric 3D representation learning and reconstruction (Müller et al. 2022; Cerkezi and Favaro 2024), pushing object-level modeling toward more realistic and scalable scenarios. In contrast, STORM adopts SAM3D as its 3D reconstruction backend (Team et al. 2025), directly constructing object-level 3D models from one or multiple reference images. Given reference RGB images and their corresponding segmentation masks, SAM3D produces high-fidelity 3D point clouds or meshes of the target object. This allows STORM to automatically obtain object-centric 3D models without manual CAD modeling or strong geometric priors, and to use these models as reliable input for downstream 6D pose estimation and tracking.
@@ 3 STORM
We propose STORM, which addresses two key limitations of prior approaches: (i) the reliance on an initial object mask and 3D model for pose estimation, and (ii) the lack of mechanisms to recover from tracking failures. STORM leverages efficient, pre-established models for object segmentation, language understanding, 3D reconstruction, and pose estimation. As a result, it provides a fully automated pipeline: given one or multiple reference images, STORM localizes and segments the target object, infers its semantic attributes, builds its 3D model, estimates its 6D pose, and performs self-correcting tracking across frames. The inference process in STORM comprises two main components: the Segmenting Object Module (SOM) and the Tracking Object Module (TOM). We now describe each component in detail.
Object Segmentation and 3D Model Generation from Reference Images
STORM robustly performs object segmentation by utilizing one or multiple provided reference images in conjunction with a novel attention-driven feature matching strategy.
@@
As shown in Figure 2 1, the reference images can be either a single or multiple clean images of the target object with a plain, single-color background, or a single or multiple masked reference images.
@@
To acquire semantic information, we feed the reference images $I_{\mathrm{reference}}$ and a short prompt $p$ into a vision–language model (gemma-3n-E4B-it) (Team 2025), which produces a concise description of the object $T$ (avoiding redundant information) and, by leveraging its prior knowledge, implicitly recovers a canonical, view-complete shape representation in the latent space. As depicted in Figure 2 2, this textual description is then encoded by CLIP into a vector $e_{t}\in\mathbf{R}^{d}$, which, together with HSFA and image registration, aligns this latent shape with the observed views and establishes consistent point–pixel correspondences.
Rather than altering the backbone’s core features, we inject $e_{t}$ into the downstream feature map $F\in\mathbf{R}^{H\times W\times C}$ by modifying its channel-wise normalization parameters. Specifically, we compute semantic offsets
@@
To bridge the distribution gap between reference images and real images and better capture the spatial information of the object, we introduce Hierarchical Spatial Fusion Attention (HSFA). As shown in Figure 2 3, HSFA simultaneously leverages the feature representations of the real and reference frames, fusing their distinct contextual information to directly obtain high-quality mask predictions. Overall, HSFA processes the input as follows:
@@
    Capturing inter-view relationships and rich spatial representations. Patch embeddings from each multi-view reference image are first processed independently using self-attention. These per-image embeddings are then concatenated across all views and passed through a joint local attention module, allowing the model to capture inter-view dependencies.
@@
    Extracting salient query representations. The query frame’s patch embeddings are processed via self-attention. To preserve the original Dino features, HSFA performs the first round of cross-attention between the query embeddings and the raw multi-view embeddings (i.e., before spatial-attention refinement). In subsequent cross-attention layers, the query frame attends to the outputs of the previous spatial-attention block, promoting faster convergence and enhancing object salience.
@@
    Composing joint representations. The Hierarchical Spatial Fusion Attention block integrates the query and multi-view embeddings, as described above. This fusion block is repeated $n$ times to iteratively refine the joint features. Finally, an upsampling convolutional block generates the mask of the target object.
@@
A detailed description of the full algorithm is provided in the Appendix. The output of the HSFA module is the segmentation mask of the target object: this mask is generated from a set of fused feature representations that jointly integrate spatial cues and contextual information from the query frame and the multi-view reference inputs. By progressively enhancing the responses of target-related regions during the hierarchical spatial fusion process, HSFA obtains a more accurate representation of the object’s shape and location, ultimately producing a high-quality target segmentation mask that can be directly used for subsequent pose estimation or tracking tasks.
@@
3D Model Generation.
As shown in Figure 2 4, Given one or multiple reference images $\{I_{\text{ref}}^{k}\}$ of the target object, which can be either raw RGB images captured against a plain, single-color background or already masked images, we obtain the corresponding segmentation masks $\{M_{\text{ref}}^{k}\}$ either by simple background color thresholding or by directly reusing the provided masks, without any additional learning or manual annotation. We further apply a depth estimator to each reference view to obtain depth maps $\{D_{\text{ref}}^{k}\}$ and use them to align the reconstruction to the correct metric scale. The masked reference images and their depths are then fed into SAM3D to reconstruct a 3D model:
$\mathcal{P}_{\text{ref}}=f_{\text{SAM3D}}\bigl(\{I_{\text{ref}}^{k},M_{\text{ref}}^{k},D_{\text{ref}}^{k}\}_{k}\bigr),$ (2)
where $\mathcal{P}_{\text{ref}}$ denotes the reconstructed, scale-consistent 3D point cloud (or mesh) of the target object. This reference 3D model is subsequently used for downstream 6D pose estimation and tracking.
@@
where $\mathrm{TP}$ is the number of true‐positive pixels of the region of interest, $\mathrm{FN}$ is the number of false‐negative pixels, and $\mathrm{FP}$ is the number of false‐positive pixels; $t\in\{0,1\}$ is the ground‐truth label for each pixel, and $p\in[0,1]$ is the predicted probability that the pixel belongs to the region of interest. The hyperparameters are set as $\alpha=0.3$, $\beta=0.7$, $\gamma=2.0$, $\lambda_{T}=1.0$, and $\lambda_{f}=0.2$. This loss formulation imposes a strong penalty on false positives, which not only effectively suppresses spurious predictions in background regions but also preserves finer discriminative ability along object boundaries. As a result, the model is able to produce compact, accurate, and detail-preserving high-quality segmentation masks that provide reliable region inputs for subsequent downstream tasks.
@@
    Initial Registration and Memory Construction. For the first tracking frame, STORM uses its segmentation module to extract an object mask. This mask, along with the generated 3D model, is passed to the FoundationPose registration module to estimate the object’s 6D pose. At the same time, a cropped RGB patch of the segmented object—referred to as a real crop—is stored in a memory pool. This crop serves as a visual reference for subsequent frames. In each new frame, the tracking module produces a pose estimate, which is compared to the memory pool to detect tracking failures. If the similarity falls below a threshold—due to rapid motion or occlusion—STORM flags the tracker as failed.
@@
    Re-initialization and Tracking Verification. When a failure is detected, STORM automatically re-initializes the tracker by SOM to generate a new segmentation mask, then re-estimating the object’s pose. A new real crop from the current frame is also added to the memory pool, ensuring that visual references stay up to date. To enable robust and generalizable comparisons, As shown in Figure 2 5, DINO is used as the feature-extraction backbone. Features from each frame are concatenated and passed through a lightweight tracking-verification subnetwork, which consists of a single self-attention layer followed by a fully connected layer and a sigmoid activation. This network outputs a binary score indicating whether tracking remains valid. The entire mechanism is designed for high efficiency, enabling real-time detection and recovery from tracking failures.
@@ 4 Experimental evaluation
    Can the 3D models generated by SAM3D provide reliable, high-quality input for downstream 6D pose estimation?
  • Q4:
@@
Zero shot
STORM (SOM, Ours) 66.1 47.1 69.1 70.8 89.L3 68.5 0.183
@@
Supervised
ZebraPoseSAT–EffNetB4 51.6 72.1 71.8 68.9 73.1 67.5 0.080
Mask R-CNN 37.5 54.4 48.9 47.1 42.9 46.2 0.100
Table 2: STORM (SOM) achieves high-quality and efficient segmentation. Comparison of zero-shot and supervised segmentation methods on five BOP datasets. mAP scores are reported on LM-O, T-LESS, TUD-L, HB, and YCB-V. Zero-shot methods, including STORM (SOM), operate without any object-specific retraining, whereas supervised baselines are fine-tuned per object category. STORM (SOM) achieves the highest mean mAP while also being highly efficient in runtime. Best values are in bold, and $\circ$ depicts the second best. SOM stands for the Segmenting Object Module in the STORM architecture.
@@
We compare STORM against the strongest publicly available systems on the BOP leaderboard. The baselines include ZebraPoseSAT-EffiNetB4 (Su et al. 2022) and Mask R-CNN (He et al. 2017), which are fine-tuned per object category, as well as NOCTIS, LDSeg, MUSE, Prisma-MPG-Complex, NIDS, SAM6D (Lin et al. 2024), CNOS, and ViewInvDet. Most baselines are distributed solely as executable packages via the BOP Challenge website without linked peer-reviewed publications, collectively referred to as the BOP Challenge Leaderboard [^1] https://bop.felk.cvut.cz/challenges/. STORM is evaluated under the same protocol for a fair comparison.
We train the SOM model on the BOP-Classic-Core training set using rendered reference images from 16 uniformly sampled spherical viewpoints, enabling HSFA to learn object-centric correspondences from these renderings (see Appendix for complete training details).
Results. Table 2 presents a comparison of segmentation methods across the BOP datasets. ZebraPoseSAT–EffNetB4 achieves a mean Average Precision (mAP) (Lin et al. 2014) of 67.567.5%, making it the strongest baseline. Our segmentation module SOM achieves a higher mAP of 68.568.5%, with an average inference time of $0.183$ seconds per image. Notably, SOM establishes a new state-of-the-art on four out of the five benchmarks—LM-O, T-LESS, HomebrewedDB, and YCB-Video—without any fine-tuning or retraining for specific object instances. These results highlight SOM’s strong generalization ability and efficiency.
Impact of 3D Model Quality on Pose Estimation
To answer Q3, we investigate whether the 3D models generated by SOM can serve as high-quality input for downstream 6D pose estimation. We evaluate STORM on the YCB-V dataset under two settings: using ground-truth 3D models as the baseline and using 3D models reconstructed from reference images as input to the downstream pose estimator. Figure 4 provides a qualitative comparison between the ground-truth CAD models and the aligned 3D models reconstructed from reference images using SAM3D. Across diverse object types, including bottles and cans, the SAM3D-generated models closely match the global geometry and fine-scale structure of the ground-truth models: object proportions, aspect ratios, and characteristic shape details (e.g., bottle necks, rims, and can lids) are faithfully preserved. Moreover, the reconstructed meshes often exhibit smoother and more regular contours along object boundaries, which can be beneficial for downstream pose estimation and rendering robustness. Only minor deviations in local surface detail and texture are observed, suggesting that SOM+SAM3D is capable of producing high-fidelity 3D geometry that is effectively indistinguishable from ground-truth CAD for the purposes of 6D pose estimation.
Results. As shown in Table 3, aligning the SAM3D-reconstructed models (SAM3D (aligned) + FP) improves pose accuracy over the unaligned variant (SAM3D (unaligned) + FP) and yields performance that is highly comparable to the ground-truth configuration (GT + FP), with only marginal differences in VSD, IoU, and ADD-S metrics. These results indicate that the 3D models produced by Aligned SAM3D are sufficiently accurate and reliable to serve as effective input for downstream 6D pose estimation.
[…]
SAM3D (unaligned) + FP 22.19 48.27 44.92
SAM3D (aligned) + FP 60.92 84.63 99.12
GT + FP 60.97 88.50 98.95
Table 3: SAM3D-generated 3D models provide reliable input for pose estimation. Mean pose metrics (VSD, IoU, ADD-S AUC) on YCB-V for SAM3D (unaligned) +Foundationpose , SAM3D (aligned) +Foundationpose, and GT+Foundationpose.
Figure 4: Comparison of 3D models reconstructed from reference images and ground-truth 3D models. The Aligned SAM3D models almost perfectly recover the underlying object structure while producing smoother and more regular contours along object boundaries.
@@
To answer Q4, we evaluate whether STORM’s tracking module (TOM) can detect when it loses track of an object—that is, whether it exhibits self-awareness of tracking failure. To this end, we synthetically generate a dataset using BOP annotations, enabling TOM to learn to distinguish between successful and failed tracking by leveraging object crops, distractors, and viewpoint variations.
@@
To answer Q5, we compare the performances of different matching strategies on the tracking-loss classification task. Using our tracking dataset, we compare STORM’s tracking module (TOM) with different configurations against a simple cosine similarity baseline. Table 5 provides the accuracy and inference time. Adding a single attention layer significantly improves accuracy from 87.5587.55% (cosine similarity) to 98.3698.36%, with only 1.31.3ms of additional latency per frame. The model without attention already achieves detection accuracy of 95.2395.23% while maintaining the fastest inference at 12.212.2ms. Adding a second attention layer does not improve the accuracy but increasing latency to 16.816.8ms, indicating that a single attention layer optimally balances accuracy and latency. TOM’s attention mechanism is crucial for robust tracking failure detection, substantially outperforming simple feature matching approaches while maintaining inference time suitable for real-time applications.